If the pressure of a gas contained in a vessel is incresed by 0.4% when heated by 1C its initial temperature must be?

Grade:11

1 Answers

Jay
6 years ago
P2=P1+P1÷250,.V=constant Tf=Ti+1,P1V/Ti=P2V/Tf,on putting values and solving,w e get.,Ti=-23℃ or 250k